# Importance of relativistic pericenter precession in identifying the presence of a third body near eccentric binaries

### v1

@@ Abstract
Many astrophysical processes can produce gravitational wave (GW) sources with significant orbital eccentricity. These binaries emit bursts of gravitational radiation during each pericenter passage. In isolated systems, the intrinsic timing of these bursts is solely determined by the properties of the binary. The presence of a nearby third body perturbs the system and alters the burst timing. Accurately modeling such perturbations therefore offers a novel approach to detecting the presence of a nearby companion. Existing timing models account for Newtonian dynamics and leading-order radiation reaction effects but neglect the higher order post-Newtonian (PN) contributions to the inner binary. In this paper, we present an improved timing model that incorporates conservative PN corrections that lead to the precession of the binary’s pericenter. We find that these PN corrections significantly impact the binary’s orbital evolution and the timing of the GW burst. In particular, 1PN precession gives rise to distinctive modulation features in the binary’s semilatus rectum and eccentricity. These modulations encode valuable information about the presence and properties of the third body, including its mass and distance. Furthermore, unmodeled 1PN effects significantly bias the tertiary’s mass and distance. Finally we assess the detectability of GW bursts from such perturbed systems and demonstrate that the inclusion of PN corrections is crucial for accurately capturing the orbital dynamics of hierarchical triples.
@@ I Introduction
Advanced LIGO (LIGO) [1] and Advanced Virgo (Virgo) [2] have detected gravitational wave (GW) signals from the mergers of more than $200$ compact binary systems during the first four observing runs [3, 4, 5, 6, 7, 8]. Many formation scenarios have been proposed for these binaries, including isolated binary evolution in the galactic field [9], dynamical interactions in stellar clusters [10, 11, 12], binaries in active galactic nuclei (AGN) [13, 14, 15, 16, 17, 18]. Current observations of binary black holes (BBHs) by the LIGO-Virgo-KAGRA (LVK) collaboration suggest that a combination of formation pathways is preferred over a single channel [19, 20, 21, 22].
@@
Orbital eccentricity of a binary is a key indicator of its formation scenario. While isolated formation channels typically produce quasi-circular binaries due to efficient circularization through radiation reaction [23], dynamically formed binaries in dense stellar environments – such as globular clusters and nuclear star clusters – can retain measurable eccentricity in the LVK frequency band [24, 25, 26, 27, 28]. Recent studies have reported weak evidence of eccentricity in a subset of events in GWTC-3 catalog [29, 30, 31, 32, 33, 34, 35, 36]. While LIGO is sensitive to BBH eccentricities $\gtrsim 0.05$ (at $10$ Hz GW frequency) [37, 38], third-generation (3G) detectors – Cosmic Explorer (CE) [39, 40] and Einstein Telescope (ET) [41, 42] – with enhanced sensitivity at low frequencies and better overall sensitivity, will be sensitive to $\sim 10$ times smaller eccentricities [43].
@@
In this paper, we mainly focus on highly eccentric binaries that emit GW radiation in the form of GW bursts. GW emission from a highly eccentric binary is suppressed for most of the orbit, except during the pericenter passage [44, 45, 46, 47, 48]. A single GW burst can have SNR $\sim 20$ in LIGO [49], whereas in the LISA band, the SNR for a single burst can reach $\mathcal{O}(100)$ [46]. The total SNR from a series of bursts roughly scales as $\sqrt{N}$ for $N$ bursts [45, 49]. As fast and accurate waveform models remain underdeveloped for highly eccentric binaries, unmodeled burst search methods such as coherent Wave Burst (cWB) [50], power stacking [51] and Bayeswave [52] could potentially detect these bursts.
@@
Various dynamical astrophysical processes can produce binaries with large eccentricities in the detector frequency band. Galactic nuclei are potential sites for forming highly eccentric binaries. The secular evolution of binaries in the vicinity of supermassive black hole (SMBH) can lead to the formation of a highly eccentric binaries through binary-single scatterings and Kozai-Lidov oscillations [53, 54, 55, 56, 57, 58]. Around $70\%$ of these BBHs can have eccentricities $\sim 0.1$ [59, 13, 15, 14, 60] and a subset of these BBHs can have eccentricity >0.9>0.9 in the LVK frequency band [61]. The presence of an AGN disc can further increase the overall merger rates in the galactic centre as well as the potential for residual eccentricity [16, 62] and also dephasing from a third object during a binary-single encounter [18] through the BH interactions with the dense gas disc.
@@
If a binary is located in the vicinity of a third body, it affects the GW burst signal in primarily two ways: (1) intrinsic changes in the inner binary’s frequency due to tidal perturbations from the third body, (2) Doppler shift due to acceleration of inner binary’s centre of mass (COM)  [^1] The Hubble expansion can also cause acceleration of the binary’s COM. However, this cosmological effect is expected to be subdominant compared to the peculiar acceleration induced by nearby astrophysical bodies [63, 64].. A notable example is a stellar mass binary orbiting a SMBH, forming a hierarchical triple system. In the Galactic centre, such triple configurations can naturally arise as the stellar mass BBH become bound to the central SMBH [65]. Tidal perturbations encode information about the intrinsic properties of the triple system, offering a novel method to probe and characterise the presence of a third body. In contrast, Doppler-induced modulations depend on the observer’s line of sight and become significant when the inner binary’s COM moves appreciably during the observation time. Both of these modulations induce a phase shift in the GW signals relative to that of an isolated binary [66, 67, 68, 69, 70, 71, 72]. Comparing the GW burst timing of a perturbed binary with that of an unperturbed binary thus provides a unique probe of the binary’s astrophysical environment in which these binaries form. Therefore, by accurately modeling the influence of a third body on the inner binary, one can detect the presence of a nearby companion and place constraints on their astrophysical environments (e.g., [73, 74]). In this work, we mainly focus on the timing modifications induced by tidal effects from a third body.
@@
Recently, Romero-Shaw etal.et\;al. [75] computed the third body effects on the timing of GW bursts from a highly eccentric binary and investigated the ability of third-generation detectors to constrain the properties of a perturbing massive third body. The timing model presented in  [75] only accounts for the Newtonian and 2.5PN order effects in the dynamics of inner binary and neglected the higher order PN effects to simplify their model. Notably, higher order PN effects are not negligible and can significantly impact the GW burst timing and the orbital evolution of the binary.
@@
More recently, Samsing etal.et\,al. [70] performed a set of N-body simulations of triple systems, with leading order conservative (Newtonian) and dissipative (2.5PN) effects. They also incorporated higher order conservative PN terms – specifically 1PN and 2PN corrections – that lead to pericenter precession of the inner binary. Their results demonstrate that neglecting these conservative PN effects in inner binary dynamics leads to an overestimation of GW phase shifts between a perturbed and a reference (unperturbed) binary. Furthermore, they found that higher order PN terms introduce additional modulation features in the binary’s orbital parameters such as semi-major axis and eccentricity.
@@
In this paper, we bridge the gap between theory and simulation by analytically incorporating conservative PN contributions to the inner binary dynamics, which induce relativistic pericenter precession of the orbit. We derive the 1PN corrections to the GW burst timing and find that these corrections significantly change the timing of the GW burst. Furthermore, 1PN corrections give rise to additional modulation features in the orbital parameters of the inner binary – features consistent with those observed in the PN simulations [70]. Moreover, in the absence of 1PN precession, the GW burst timing strongly depends on the initial argument of pericenter. Once pericenter precession is accounted for, however, this dependence becomes substantially weaker. These precession-induced effects have important implications for detecting and characterizing third body companions in eccentric binary systems. Therefore, PN corrections are crucial for accurately modeling the GW burst timing and orbital evolution of the binary.
@@ II.1 Osculating orbit formalism
Figure 1 shows the geometric illustration of a triple system. A test particle of reduced mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ moves in the gravitational potential of a total mass $m=m_{1}+m_{2}$, located at the centre of mass of the inner binary. We denote the true anomaly of the inner orbit by $V=\phi-\omega$, where $\phi$ is the orbital phase and $\omega$ is the angle of the pericenter. The true anomaly of the outer orbit is represented by angle $V_{3}$. The orbital eccentricity is denoted by $e$, the semi-major axis by $a$ and $p=a(1-e^{2})$ represents the semi-latus rectum of the inner binary. The orientation of the source with respect to the detector is represented by inclination angle $\iota$.
Figure 1: A geometric representation of a triple system. The inner binary follows an eccentric orbit with reduced mass $\mu$ orbiting the total mass $m$ at a separation $r$. The third body of mass $m_{3}$ is located at a much larger distance $R\gg r$ and moves on a circular outer orbit. The angle $V$ denotes the true anomaly of the inner orbit and $\omega$ is the longitude of pericenter. The true anomaly of the outer orbit is $V_{3}$. Both orbits are assumed to be coplanar. The inclination angle $\iota$ describes the orientation of the system relative to the detector.
The system is an example of a perturbed Kepler problem. To solve it, we employ the method of osculating orbits [76]. Due to the presence of a small perturbing force, the otherwise constant orbital elements of the binary evolve slowly over time. As a result, the dynamics of the binary can be approximated by a series of instantaneous Keplerian orbits whose parameters evolve secularly due to the perturbing force. The relative acceleration between two bodies $(\bm{a}=\bm{a}_{1}-\bm{a}_{2})$ can be written in the form
@@
where $\bm{n}=\bm{r}/r$, $|\bm{r}|=r$ and $\bm{f}$ is a perturbing force per unit mass and the radial distance to the test particle is given by
@@
$\bm{f}=\mathcal{R}\bm{n}+\mathcal{S}\bm{\lambda}+\mathcal{W}\bm{e_{z}}\,$, (3)
@@
The following equations describe the vectorial basis
@@
where $\bm{N}=\bm{R}/R$. The orbital evolution is governed by the following system of osculating equations [76]
@@ II.2.1 Stability
For the coplanar and circular outer orbit, the stability criterion for a triple system is given by [77, 78]
$\frac{R}{m_{3}}\gtrsim 2.8\frac{(p/m)}{(m_{3}/m)^{3/5}(1-e^{2})}\,$, (8)
This criterion ensures long-term dynamical stability by preventing the third body from disrupting the inner binary.
@@ II.2.2 Chirping outer binary
We neglect the PN corrections and radiation reaction effects on the outer orbit. This assumption is justified because we assume that the third body is located at a distance much larger than the binary separation i.e., $R\gg r$. To leading order, the radius of a circular outer orbit evolves according to the following expression:
@@
where $\eta_{\rm out}=(m_{3}m)/(m+m_{3})^{2}$. To quantify the effect of radiation reaction on the outer orbit, we require that the radius of the outer orbit does not evolve significantly during the observation time i.e., $\delta R/R_{0}\lesssim 1\%$ [79].
$\frac{R}{m_{3}}\gtrsim\left(\frac{m}{m_{3}}\right)\left[1280\;\frac{\eta_{\rm out}}{m}\left(1+\frac{m_{3}}{m}\right)^{3}P_{\rm obs}\right]^{1/4}.$ (10)
@@ II.2.3 Orbital timescales
Figure 2: The allowed values of tertiary distance $R$ and mass $m_{3}$ for various physical constraints discussed in Sec. II.2. Different colors represent different physical conditions. The solid lines denote $p=30m$ and dashed lines show $p=200m$. We choose an equal mass binary with $m=30M_{\odot}$, and $e=0.99$. The horizontal line represents the radius of the innermost stable circular orbit for a maximally spinning black hole $(R_{\rm min}=m_{3})$. The arrows denote the region of allowed parameters.
@@
In eccentric systems, GW emission occurs not only at the fundamental orbital frequency but also at its integer harmonics. For highly eccentric binaries, the radiation is dominated by higher harmonics, leading to a spectral shift toward higher GW frequencies [80]. The peak GW frequency, corresponding to the harmonic with the maximum power, can be estimated using the following expression:  [^2] Note that this expression becomes inaccurate for low eccentricities. For low eccentricity binaries, Ref. [81] provided an improved numerical fit for $f_{\rm peak,GW}$. Since we consider only highly eccentric binaries in this study, Eq. (12) is good enough for our purpose. [82]
@@
Here $n_{\rm peak}$ is the peak harmonic number and $f_{\rm orb}$ is the orbital frequency. Most of the sources considered in this work emit GWs with peak frequencies that lie within the sensitivity band of ground-based detectors such as LIGO, CE, and ET. For example, an equal mass binary with $m=30M_{\odot}$, $p=200m$, $e_{0}=0.99$ will have a peak frequency of $\sim 2$Hz –lower cut-off frequency of Einstein Telescope.

### v2

@@ Abstract
Many astrophysical processes can produce gravitational wave (GW) sources with significant orbital eccentricity. These binaries emit bursts of gravitational radiation during each pericenter passage. In isolated systems, the intrinsic timing of these bursts is solely determined by the properties of the binary. The presence of a nearby third body perturbs the system and alters the burst timing. Accurately modeling such perturbations therefore offers a novel approach to detecting the presence of a nearby companion. Existing timing models account for Newtonian dynamics and leading order radiation reaction effects but neglect the higher order post-Newtonian (PN) contributions to the inner binary. In this paper, we present an improved timing model that incorporates conservative PN corrections that lead to the precession of the binary’s pericenter. We find that these PN corrections significantly impact the binary’s orbital evolution and the timing of the GW burst. In particular, 1PN precession gives rise to distinctive modulation features in the binary’s semilatus rectum and eccentricity. These modulations encode valuable information about the presence and properties of the third body, including its mass and distance. Furthermore, unmodeled 1PN effects significantly bias the tertiary’s mass and distance. Finally, we assess the detectability of GW bursts from such perturbed systems and demonstrate that the inclusion of PN corrections is crucial for accurately capturing the orbital dynamics of hierarchical triples.
@@ I Introduction
Advanced LIGO (LIGO) [LIGOScientific:2014pky] and Advanced Virgo (Virgo) [VIRGO:2014yos] have detected gravitational wave (GW) signals from the mergers of more than $200$ compact binary systems during the first four observing runs [LIGOScientific:2020ibl, KAGRA:2021duu, KAGRA:2021vkt, Nitz:2021zwj, LIGOScientific:2025rsn]. Many formation scenarios have been proposed for these binaries, including isolated binary evolution in the galactic field [Bavera:2020uch], dynamical interactions in stellar clusters [Chatterjee:2016thb, Mandel:2018hfr, DiCarlo:2019pmf], and binaries in active galactic nuclei (AGN) [Tagawa:2019osr, Samsing:2020tda, Tagawa:2020jnc, Rowan:2022ehz, Whitehead:2023hmh, Rowan:2025xxb]. Current observations of binary black holes (BBHs) by the LIGO-Virgo-KAGRA (LVK) collaboration suggest that a combination of formation pathways is preferred over a single channel [LIGOScientific:2020kqk, Bouffanais:2021wcr, PhysRevD.103.023026, Zevin:2020gbd].
@@
Orbital eccentricity of a binary is a key indicator of its formation scenario. While isolated formation channels typically produce quasi-circular binaries due to efficient circularization through radiation reaction [PhysRev.136.B1224], dynamically formed binaries in dense stellar environments – such as globular clusters and nuclear star clusters – can retain measurable eccentricity in the LVK frequency band [Samsing:2013kua, PhysRevD.97.103014, Rasskazov:2019gjw, Zevin:2021rtf, DallAmico:2023neb]. Recent studies have reported weak evidence of eccentricity in a subset of events in GWTC-3 catalog [Romero-Shaw:2020thy, Gayathri:2020coq, Romero-Shaw:2021ual, OShea:2021faf, Gupte:2024jfe, Morras:2025xfu, Jan:2025fps, Kacanja:2025kpr]. While LIGO is sensitive to BBH eccentricities $\gtrsim 0.05$ (at $10$ Hz GW frequency) [PhysRevD.98.083028, Favata:2021vhw], third-generation (3G) detectors – Cosmic Explorer (CE) [LIGOScientific:2016wof, Reitze:2019iox] and Einstein Telescope (ET) [Punturo:2010zz, Sathyaprakash:2011bh] – with enhanced sensitivity at low frequencies and better overall sensitivity, will be sensitive to $\sim 10$ times smaller eccentricities [Saini:2023wdk].
@@
In this paper, we mainly focus on highly eccentric binaries that emit GW radiation in the form of GW bursts. GW emission from a highly eccentric binary is suppressed for most of the orbit, except during the pericenter passage [PhysRevD.90.103001, Loutrel:2017fgu, Xuan:2023azh, PhysRevD.85.123005, PhysRevD.110.023020]. A single GW burst can have SNR $\sim 20$ in LIGO [Loutrel:2020jfx], whereas in the LISA band, the SNR for a single burst can reach $\mathcal{O}(100)$ [Xuan:2023azh]. The total SNR from a series of bursts roughly scales as $\sqrt{N}$ for $N$ bursts [Loutrel:2017fgu, Loutrel:2020jfx]. As fast and accurate waveform models remain underdeveloped for highly eccentric binaries, unmodeled burst search methods such as coherent Wave Burst [Klimenko:2008fu], power stacking [Tai:2014bfa], and Bayeswave [Cornish:2014kda] could potentially detect these bursts.
@@
Various dynamical astrophysical processes can produce binaries with large eccentricities in the detector frequency band. Galactic nuclei are potential sites for forming highly eccentric binaries. The secular evolution of binaries in the vicinity of a supermassive black hole (SMBH) can lead to the formation of a highly eccentric binaries through binary-single scatterings and Kozai-Lidov oscillations [2010ApJ...713...90A, 2015ApJ...799..118P, 2013MNRAS.431.2155N, Hoang:2017fvh, 2022ApJ...925..178H, Su_2025_kozai, Camilloni:2023xvf]. Around $70\%$ of these BBHs can have eccentricities $\sim 0.1$ [Takatsy:2018euo, Tagawa:2019osr, 2019MNRAS.488...47F, Tagawa:2020jnc, Samsing:2020tda, Gondan:2020svr], and a subset of these BBHs can have eccentricity >0.9>0.9 in the LVK frequency band [OLeary:2008myb]. The presence of an AGN disc can further increase the overall merger rates in the galactic center as well as the potential for residual eccentricity [Rowan:2022ehz, Calcino2024, 2025arXiv250803637V, 2025ApJ...982L..13G] and also dephasing from a third object during a binary-single encounter [Rowan:2025xxb] through the BH interactions with the dense gas disc.
@@
If a binary is located in the vicinity of a third body, it affects the GW burst signal in primarily two ways: (1) intrinsic changes in the inner binary’s frequency due to tidal perturbations from the third body, (2) Doppler shift due to acceleration of inner binary’s center of mass (COM).  [^1] The Hubble expansion can also cause acceleration of the binary’s COM. However, this cosmological effect is expected to be subdominant compared to the peculiar acceleration induced by nearby astrophysical bodies [PhysRevD.83.044030, PhysRevD.95.044029]. A notable example is a stellar mass binary orbiting a SMBH, forming a hierarchical triple system. In the Galactic Center, such triple configurations can naturally arise as the stellar mass BBH become bound to the central SMBH [Grishin:2021hcp]. Tidal perturbations encode information about the intrinsic properties of the triple system, offering a novel method to probe and characterize the presence of a third body. In contrast, Doppler-induced modulations depend on the observer’s line of sight and become significant when the inner binary’s COM moves appreciably during the observation time. Both of these modulations induce a phase shift in the GW signals relative to that of an isolated binary [Meiron:2016ipr, Inayoshi:2017hgw, PhysRevD.105.024017, Vijaykumar:2023tjg, Tiwari:2023cpa, Samsing:2024syt, Hendriks:2024zbu, Hendriks:2024gpp, l542-s32g]. Comparing the GW burst timing of a perturbed binary with that of an unperturbed binary thus provides a unique probe of the binary’s astrophysical environment in which these binaries form. Therefore, by accurately modeling the influence of a third body on the inner binary, one can detect the presence of a nearby companion and place constraints on their astrophysical environments [e.g., [Yunes:2010sm, Laeuger:2023qyz]]. In this work, we mainly focus on the timing modifications induced by tidal effects from a third body.
@@
Recently, Romero-Shaw etal.et\;al. [Isobel] computed the third body effects on the timing of GW bursts from a highly eccentric binary and investigated the ability of third-generation detectors to constrain the properties of a perturbing massive third body. The timing model presented in  [Isobel] only accounts for the Newtonian and 2.5PN order effects in the dynamics of the inner binary and neglected the higher order post-Newtonian (PN) effects to simplify their model. Notably, higher order PN effects are not negligible and can significantly impact the GW burst timing and the orbital evolution of the binary.
@@
More recently, Samsing etal.et\,al. [Samsing:2024syt] performed a set of N-body simulations of triple systems, with leading order conservative (Newtonian) and dissipative (2.5PN) effects. They also incorporated higher order conservative PN terms – specifically 1PN and 2PN corrections – that lead to pericenter precession of the inner binary. Their results demonstrate that neglecting these conservative PN effects in inner binary dynamics leads to an overestimation of GW phase shifts between a perturbed and a reference (unperturbed) binary. Furthermore, they found that higher order PN terms introduce additional modulation features in the binary’s orbital parameters such as semi-major axis and eccentricity.
@@
In this paper, we bridge the gap between theory and simulation by analytically incorporating conservative PN contributions to the inner binary dynamics, which induce relativistic pericenter precession of the orbit. We derive the 1PN corrections to the GW burst timing and find that these corrections significantly change the timing of the GW burst. Furthermore, 1PN corrections give rise to additional modulation features in the orbital parameters of the inner binary – features consistent with those observed in the PN simulations [Samsing:2024syt]. Moreover, in the absence of 1PN precession, the GW burst timing strongly depends on the initial argument of pericenter. Once pericenter precession is accounted for, however, this dependence becomes substantially weaker. These precession-induced effects have important implications for detecting and characterizing third body companions in eccentric binary systems. Therefore, PN corrections are crucial for accurately modeling the GW burst timing and orbital evolution of the binary.
@@ II.1 Osculating orbit formalism
Figure 1 shows the geometric illustration of a triple system. A test particle of reduced mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ moves in the gravitational potential of a total mass $m=m_{1}+m_{2}$, located at the center of mass of the inner binary. We denote the true anomaly of the inner orbit by $V=\phi-\omega$, where $\phi$ is the orbital phase and $\omega$ is the angle of the pericenter. The true anomaly of the outer orbit is represented by angle $V_{3}$. The orbital eccentricity is denoted by $e$, the semi-major axis by $a$, and $p=a(1-e^{2})$ represents the semi-latus rectum of the inner binary. The orientation of the source with respect to the detector is represented by inclination angle $\iota_{3}$. We assume $\iota_{3}=0$ throughout, except in Sec. IV.3.
Figure 1: Geometric representation of a triple system. The inner binary follows an eccentric orbit with reduced mass $\mu$ orbiting the total mass $m$ at a separation $r$. The third body of mass $m_{3}$ is located at a much larger distance $R\gg r$ and moves on a circular outer orbit. The angle $V$ denotes the true anomaly of the inner orbit, and $\omega$ is the longitude of pericenter. The true anomaly of the outer orbit is $V_{3}$. Both orbits are assumed to be coplanar. The inclination angle $\iota_{3}$ describes the orientation of the system relative to the detector.
The system is an example of a perturbed Kepler problem. To solve it, we employ the method of osculating orbits [Poisson_Will_2014]  [^2] See also [2025arXiv250806245G] for a unified perturbed Kepler framework incorporating a large class of deviations from Keplerian binary due to environmental effects.. Due to the presence of a small perturbing force, the otherwise constant orbital elements of the binary evolve slowly over time. As a result, the dynamics of the binary can be approximated by a series of instantaneous Keplerian orbits whose parameters evolve secularly due to the perturbing force. The relative acceleration between two bodies $(\bm{a}=\bm{a}_{1}-\bm{a}_{2})$ can be written in the form
@@
where $\bm{n}=\bm{r}/r$ is a unit vector that points from body 2 to body 1, $|\bm{r}|=r$ is the separation between bodies, $\bm{f}$ is a perturbing force per unit mass, and the radial distance to the test particle is given by
@@
$\bm{f}=\mathcal{R}\bm{n}+\mathcal{S}\bm{\lambda}+\mathcal{W}\bm{e_{z}}\,.$ (3)
@@
The following equations describe the vectorial basis:
@@
where $\bm{N}=\bm{R}/R$. The orbital evolution is governed by the following system of osculating equations [Poisson_Will_2014]:
@@ II.2.1 Stability
For the coplanar and circular outer orbit, the stability criterion for a triple system is given by [2001MNRAS.321..398M, Blaes:2002cs]
$\frac{R}{m_{3}}\gtrsim 2.8\frac{(p/m)}{(m_{3}/m)^{3/5}(1-e^{2})}\,.$ (8)
This criterion ensures long-term dynamical stability by preventing the third body from disrupting the inner binary. Note that mutually inclined orbits are expected to be more stable [Blaes:2002cs], so Eq. (8) provides a more conservative stability limit. [See also [2022MNRAS.516.4146V, 2022PASA...39...62T] for outer eccentric orbits and various mass ratios.]
@@ II.2.2 Chirping outer binary
We neglect the PN corrections and radiation reaction effects on the outer orbit. This assumption is justified because we assume that the third body is located at a distance much larger than the binary separation, i.e., $R\gg r$. To leading order, the radius of a circular outer orbit shrinks according to the following expression [PhysRev.136.B1224]:
@@
where $\eta_{\rm out}=(m_{3}m)/(m+m_{3})^{2}$. To quantify the effect of radiation reaction on the outer orbit, we require that the radius of the outer orbit does not evolve significantly during the observation time, i.e., $\delta R/R_{0}\lesssim 1\%$ [Chandramouli:2021kts].
$\frac{R}{m_{3}}\gtrsim\left(\frac{m}{m_{3}}\right)\left[1280\;\frac{\eta_{\rm out}}{m}\left(1+\frac{m_{3}}{m}\right)^{3}P_{\rm obs}\right]^{1/4},$ (10)
@@ II.2.3 Orbital timescales
Figure 2: Allowed values of tertiary distance $R$ and mass $m_{3}$ for various physical constraints discussed in Sec. II.2. Different colors represent different physical conditions. Solid lines denote $p=30m$ and dashed lines show $p=200m$. We choose an equal mass binary with $m=30M_{\odot}$ and $e=0.99$. The horizontal line represents the radius of the innermost stable circular orbit for a maximally spinning black hole $(R_{\rm min}=m_{3})$. The arrows denote the region of allowed parameters.
@@
In eccentric systems, GW emission occurs not only at the fundamental orbital frequency but also at its integer harmonics. For highly eccentric binaries, the radiation is dominated by higher harmonics, leading to a spectral shift toward higher GW frequencies [PhysRev.131.435]. The peak GW frequency, corresponding to the harmonic with the maximum power, can be estimated using the following expression  [^3] Note that this expression becomes inaccurate for low eccentricities. For low eccentricity binaries, Ref. [Hamers:2021eir] provided an improved numerical fit for $f_{\rm peak,GW}$. Since we consider only highly eccentric binaries in this study, Eq. (12) is good enough for our purpose. [Wen:2002km]:
@@
Here $n_{\rm peak}$ is the peak harmonic number and $f_{\rm orb}$ is the orbital frequency. Most of the sources considered in this work emit GWs with peak frequencies that lie within the sensitivity band of ground-based detectors such as LIGO, CE, and ET. For example, an equal mass binary with $m=30M_{\odot}$, $p=200m$, $e_{0}=0.99$ will have a peak frequency of $\sim 2$Hz, lower cut-off frequency of Einstein Telescope.